OpenStudy (anonymous):

I don't understand.... Given the functions f(x)= x+2 and g(x) = x^2, what values of x result in f(g(x))=g(f(x))?

OpenStudy (tkhunny):

Have you considered creating the two functions and solving the resultant equation?

OpenStudy (shamil98):

^

OpenStudy (shamil98):

f(g(x)) = f o g g(f(x)) = g o f so f(g(x)) = x^2 + 2 g(f(x)) = (x+2)^2 now just set these two equal to each other and solve.. x^2 + 2 = (x+2)^2